OpenStudy (anonymous):

Ok, so I've got this math homework and it's a bit confusing. Can you please help? A 9.5m length of timber is used to make a rectangular frame who's dimensions are in the ratio of 3:2. Find the length and width of the frame. Please don't answer, just explain how! Thanks

OpenStudy (anonymous):

if x and y are taken as length and width, x:y=3:2, x/y=3/2,x=3y/2 2(x+y)=9.5 2(3y/2+y)=9.5 5y=9.5 y=1.9 x=2.85

OpenStudy (anonymous):

rectangle is made by a length of 9.5, which is the perimeter